A 10g bullet traveling at 400 m/s strikes a 10 kg,
Answer #1

here,

mass of the bullet , m = 0.01 kg

speed , u = 400 m/s

mass , M = 10 kg

width of door , r = 1 m

let the angular speed of the door after the impact be w

using conservation of angular momentum about the hinge

m * v * r = ( M + m) * r^2 * w

0.01 * 400 = ( 10 + 0.01) * 1 * w

w = 0.4 rad/s

the angular speed of the door is 0.4 rad/s